OpenStudy (anonymous):

4. Find a third degree polynomial equation with rational coefficients that has roots -4 and 6+i (1 Point) A. x^3-8x^2-11x+148=0 B. x^3-8x^2-12x^2+37x=0 C.x^3-12x^2+37x=0 D.x^3-8x^2-11x=0

OpenStudy (anonymous):

4. Find a third degree polynomial equation with rational coefficients that has roots -4 and 6+i (1 Point) A. (1 point) x^3-8x^2-11x+148=0 B. (0 points) x^3-8x^2-12x^2+37x=0 C. (0 points) x^3-12x^2+37x=0 D. (0 points) x^3-8x^2-11x=0